# Towards Intelligent Battery Management via A Five-Tier Digital Twin Framework

### v1

@@ III-A3 AI Engine Module
From a system-level perspective, the AI engine drives autonomous operation by turning forecasts and optimizations into direct control actions. It utilized continual learning to adapt operational policies as usage conditions vary, and employs fault-tolerant mechanisms that keep the system functional when components fail. Before new decisions are enacted, the engine runs fast predictive simulations to conduct risk detections such as thermal runaway or internal short-circuit. To validate these learned policies work effectively on real battery systems, the AI engine is coupled with HIL platforms (e.g., OPAL-RT), which execute virtual commands against real controllers and components. In this way, optimized policies are validated under realistic constraints before deployment, ensuring that the digital twin can operate as a self-governing battery manager across its lifecycle. Overall, these capabilities establish the AI engine as the intelligence core of the digital twin, as illustrated in Figure 3.
Figure 3: Integrated Technical Architecture of the Battery Digital Twin Platform. The comprehensive technical architecture of our digital twin platform includes multi-physics modeling tools (e.g., SPM and DFN models), advanced simulation platforms, and physics-informed AI methodologies such as PhysicsNeMo. The layered structure enables systematic data integration, accurate predictive simulations, and effective prescriptive strategies, thus supporting proactive battery health management throughout its lifecycle.
III-B Potential Applications
The capabilities of our five-tier digital twin framework enable transformative applications across battery management, through two primary domains that collectively address intelligent battery operations.
III-B1 Predictive Battery Health Management
The most immediate value of the proposed battery digital twin lies in its ability to transform health prediction into a continuous and adaptive process. Instead of relying on sparse snapshots, SOH and RUL are evaluated in real time, adapting to variations in usage and environment. This approach improves predictive accuracy and produces risk-aware outputs expressed as probability distributions rather than deterministic forecasts, enabling operators to plan maintenance proactively with quantified confidence. Such predictive intelligence ensures that maintenance actions are both timely and cost-effective, reducing unexpected downtime and improving overall system reliability.
@@ III-B1 Predictive Battery Health Management
III-B2 Battery Repurposing and Value Maximization
@@
Beyond health prediction, the digital twin enables actionable strategies to maximize battery value across its entire lifecycle. During operation, forecasts are translated into adaptive charging protocols that shorten charging duration while keeping temperature rise and energy loss within safe limits, and intelligent thermal management that dynamically adjusts airflow and coolant flow according to predicted heat loads. These strategies not only improve energy efficiency but also mitigate thermal stress that accelerates degradation. At the system level, the twin orchestrates load distribution across modules with different health states. Instead of uniform current sharing, it redistributes demand away from overstressed or degraded packs, thereby protecting weaker components while extracting maximum capacity from healthier ones. This capability is particularly valuable in grid-scale storage plants integrating batteries of mixed chemistries and vintages, where conventional rule-based control often leads to underutilization of robust modules and premature aging of already stressed ones [12].
The framework further supports autonomous lifecycle management by combining health forecasts with economic reasoning. Instead of fixed service intervals, it continuously evaluates the cost–benefit trade-offs of interventions such as rebalancing or component replacement, ensuring maintenance actions are both timely and economically justified. Finally, by preserving complete degradation histories, the digital twin provides accurate assessments of residual capacity at end-of-life. This enables reliable decisions for second-life deployment in stationary storage or recycling, ensuring safe, efficient, and sustainable reuse while maximizing residual value.
IV Case Study: Integrated Multi-physics Calibration and Uncertainty-quantified SOH Prediction
To validate the effectiveness of the proposed framework, we present a case study on multi-physics calibration of cell voltage and temperature, and uncertainty-quantified SOH prediction for battery operation.
IV-A Problem Settings
This case study demonstrates the practical application and effectiveness of our proposed digital twin framework for intelligent battery management using the XJTU battery dataset [13]. The dataset consists of run-to-end of life experiments performed on 55 cylindrical 18650 Nickel-Cobalt-Manganese (NCM) lithium-ion batteries ($\mathrm{LiNi_{0.5}Co_{0.2}Mn_{0.3}O_{2}}$), each with a nominal capacity of $2000\,\mathrm{mAh}$ and nominal voltage of $3.6\,\mathrm{V}$. The experiments are carried out using six different charge-discharge regimes designed to emulate various realistic operational conditions, ranging from constant C-rate cycling and variable discharge conditions to randomized and specialized satellite battery usage patterns.
In our study, we combine physical modeling with data-driven models to address two core functions of practical battery management. First of all, we employ a DFN electrochemical model combined with a lumped thermal model to simulate the battery operation. Accurate calibration of this multi-physics model is the key to reliably simulating the voltage and temperature behavior of batteries under diverse cycling conditions, as represented in the dataset. Calibrating these models is fundamental, as precise physical simulations constitute the cornerstone of an effective digital twin for advanced battery management, as we addressed before. Second, utilizing the calibrated simulation data sets up the foundation to develop an advanced predictive model to accurately predict SOH while simultaneously quantifying the uncertainty. The integration of uncertainty quantification addresses a critical gap in existing deterministic prediction methodologies, enabling risk-informed operational decision-making for improved battery lifecycle management.
IV-B Model Architecture
Our integrated battery digital twin combines electrochemical-thermal simulations, Physics-informed Neural Networks (PINNs), and uncertainty quantification through Deep Autoencoding Gaussian Mixture Models (DAGMM). This approach enables accurate and physically consistent predictions of battery SOH and robust quantification of prediction uncertainty.
IV-B1 Multi-physics Modeling and Calibration
@@
We integrate the SPM and DFN electrochemical models with a lumped thermal model using PyBaMM as the model source platform [14]. The SPM offers computationally efficient simulations suitable for rapid, real-time predictions, while the DFN model provides detailed modeling of internal lithium-ion transport processes and reaction kinetics. Lumped thermal model is incorporated with the DFN model to ensure an accurate representation of thermal dynamics that effectively captures dynamic battery behavior under various operational conditions.
@@
To ensure the fidelity of our simulation models with respect to actual battery performance, we further implement a rigorous Bayesian optimization procedure for parameter calibration. This optimization systematically adjusts over 15 model parameters, such as diffusivities, reaction rate, conductivities, electrode geometries, and thermal conductivities, by minimizing discrepancies between simulated outputs and experimentally measured voltage and temperature data. Through iterative refinement using Bayesian optimization methods, our calibrated multi-physics models attain strong consistency with observed battery responses, ensuring reliable simulation outcomes.
IV-B2 Physics-informed Neural Network
@@
Leveraging NVIDIA PhysicsNeMo, we construct PINNs that combine deep learning approaches with PyBaMM simulations. These PINNs are trained using data from our calibrated multi-physics simulations, incorporating a physics-informed loss function as the learning bias that integrates both data-driven accuracy terms and physical constraints. Specifically, the loss function penalizes deviations from fundamental battery physics and thermal equilibrium. By enforcing these constraints via automatic differentiation at strategically selected points within the battery simulation domain, this proposed PINN approach effectively captures the complex battery dynamics while maintaining adherence to underlying physical laws.
IV-B3 Uncertainty Quantification via DAGMM
In addition to these approaches mentioned above, our predictive model further incorporates DAGMM for uncertainty quantification. DAGMM leverages an autoencoder network to transform operational data into latent representations, subsequently modeled by a Gaussian Mixture Model (GMM). The uncertainty metric is obtained through an energy score, evaluating latent representations’ likelihood within the learned GMM. This energy-based score identifies distributional anomalies indicative of prediction uncertainty.
IV-C Evaluation Results
@@
Rigorous evaluations are given to demonstrate the superior predictive accuracy and reliability of the proposed digital twin framework through multi-physics model calibration and uncertainty quantifications.
@@
Figure 4: Multi-physics calibration results against the XJTU battery dataset after Bayesian optimization. (a) Voltage prediction under 1C, 2C, and 3C discharge rates. (b) Temperature prediction under the same operating conditions.
@@
For the multi-physics calibration, the electrochemical-thermal model demonstrated good estimation accuracy compared with the XJTU battery dataset with the use of Bayesian optimization as shown in Figure 4. For voltage prediction, the calibrated model achieved an MAPE of 0.92% (1C), 1.06% (2C), and 1.57% (3C). For temperature prediction, the MAPE reached 0.07% (1C), 0.18% (2C), and 0.39% (3C). These results validate the effectiveness of the multi-physics model that can accurately capture both potential and thermal responses of batteries under different operational regimes.
@@
Building upon the calibrated simulations, we can further assess the predictive performance of our PINN-based SOH prediction model, including its capability to quantify prediction uncertainty. As shown in Figure 5, our model demonstrated robust predictive accuracy in SOH, achieving mean absolute percentage errors (MAPE) consistently below 3%. Most importantly, the integrated uncertainty quantification via the DAGMM provided reliable indications of prediction confidence. As shown in Figure 6, the energy-based uncertainty scores exhibited a strong positive correlation with the actual prediction errors, effectively highlighting scenarios where predictions were less reliable due to data distribution shifts.
@@
These comprehensive evaluation outcomes validate that our integrated digital twin framework not only achieves precise physical modeling and accurate health predictions but also effectively quantifies uncertainties, significantly enhancing the robustness and reliability of battery management decisions.
V Future Directions
@@
Based on the initial results from the case study mentioned above, we discuss several future directions that can be explored to further enhance the intelligent battery management using this digital twin architecture.
@@
Foundation Models for Battery Intelligence: Large Language Models (LLMs) are foundation models trained on massive text corpora using transformer architectures, enabling diverse tasks such as reasoning, summarization, and question answering without task-specific supervision. Beyond language, they serve as general-purpose engines for knowledge representation and code generation across scientific domains. Building on these capabilities, LLMs adapted on battery-related literature present promising opportunities for digital twins through automated knowledge extraction and model generation [15]. Future implementations envision specialized battery foundation models that automatically design PINN architectures for specific chemistries, synthesize insights from vast research to identify degradation mechanisms, recommend experimental protocols, and support conversational interfaces for intuitive querying of battery states and explanations of complex degradation phenomena.
@@
Blockchain-based Battery Passport Systems: The implementation of blockchain in battery lifecycle management offers transformative potential for future battery passport systems, enabling complete traceability and transparency. Distributed ledger technologies will record manufacturing data, operational history, maintenance, and performance metrics as immutable entries, ensuring reliable information sharing across manufacturers, operators, and recyclers. Smart contracts could automate management decisions such as maintenance scheduling and end-of-life processing, reducing fraud in condition reporting. Blockchain-based passports will further support circular economy initiatives by providing verified health data for second-life applications, allowing batteries retired from vehicles to be reused in stationary storage with confidence.
@@
Differentiable Simulation for Optimal Control: The development of fully differentiable battery simulation frameworks enables gradient-based optimization of battery operational strategies directly through physics-based models, eliminating the need for computationally expensive reinforcement learning approaches. Future implementations could leverage automatic differentiation through simulation platforms to enable direct optimization of charging protocols, thermal management strategies, and load balancing decisions with respect to battery health and performance objectives. Differentiable simulation supports real-time model predictive control that optimizes battery operation while accounting for complex multi-physics interactions and degradation constraints. This approach provides enhanced transparency compared to black-box optimization methods while enabling principled handling of operational constraints through physics-based penalty terms in the optimization objective.
VI Conclusion
@@
In this article, we presented an intelligent battery management concept based on a five-tier digital twin framework for autonomous operations. We first highlighted current BMS challenges and the need for advanced digital twin approaches. Then, we introduced the proposed framework, progressing from geometric modeling to autonomous operation, supported by PIML and uncertainty quantification. Its applications span predictive health management, operational optimization, and lifecycle control. Finally, a case study validated our design, showing sub-1% voltage/temperature errors and robust SOH predictions with MAPE below 3%, demonstrating the framework’s potential for interpretable modeling and reliable decision-making.

### v2

@@ Abstract
From a system-level perspective, the AI engine drives autonomous operation by turning forecasts and optimizations into direct control actions. It utilized continual learning to adapt operational policies as usage conditions vary, and employs fault-tolerant mechanisms that keep the system functional when components fail. Before new decisions are enacted, the engine runs fast predictive simulations to conduct risk detections such as thermal runaway or internal short-circuit. To validate these learned policies work effectively on real battery systems, the AI engine is coupled with HIL platforms (e.g., OPAL-RT), which execute virtual commands against real controllers and components. In this way, optimized policies are validated under realistic constraints before deployment, ensuring that the digital twin can operate as a self-governing battery manager across its lifecycle.
Figure 3: Technical architecture of the five-tier intelligent digital twin system for battery management. The comprehensive technical architecture includes multi-physics modeling tools, advanced simulation platforms, and physics-informed AI methodologies such as PhysicsNeMo. The layered structure enables systematic data integration, accurate predictive simulations, and effective prescriptive strategies, thus supporting proactive battery health management throughout its lifecycle.
Practical Applications of The Proposed Intelligent System
The capabilities of our intelligent digital twin system enable transformative applications across battery management, through two primary domains that collectively address intelligent battery operations.
Predictive Battery Health Management
The most immediate value of the proposed battery digital twin lies in its ability to transform health prediction into a continuous and adaptive process. Instead of relying on sparse snapshots, SOH and RUL are evaluated in real time, adapting to variations in usage and environment. This approach improves predictive accuracy and produces risk-aware outputs expressed as probability distributions rather than deterministic forecasts, enabling operators to plan maintenance proactively with quantified confidence [gama2022data]. Such predictive intelligence ensures that maintenance actions are both timely and cost-effective, reducing unexpected downtime and improving overall system reliability.
@@
Battery Repurposing and Value Maximization
@@
Beyond health prediction, the digital twin enables actionable strategies to maximize battery value across its entire lifecycle. During operation, forecasts are translated into adaptive charging protocols that shorten charging duration while keeping temperature rise and energy loss within safe limits, and intelligent thermal management that dynamically adjusts airflow and coolant flow according to predicted heat loads. These strategies not only improve energy efficiency but also mitigate thermal stress that accelerates degradation. At the system level, the twin orchestrates load distribution across modules with different health states. Instead of uniform current sharing, it redistributes demand away from overstressed or degraded packs, thereby protecting weaker components while extracting maximum capacity from healthier ones. This capability is particularly valuable in grid-scale storage plants integrating batteries of mixed chemistries and vintages, where conventional rule-based control often leads to underutilization of robust modules and premature aging of already stressed ones [koltermann2024improved]. Finally, by preserving complete degradation histories, the digital twin provides accurate assessments of residual capacity at end-of-life. This enables reliable decisions for second-life deployment in stationary storage or recycling, ensuring safe, efficient, and sustainable reuse while maximizing residual value.
System Implementation and Evaluation
In this section, we implement the proposed five-tier intelligent digital twin system based on NVIDIA ecosystem, focusing on model calibration and health prediction for practical battery management.
Problem Settings
XJTU battery dataset [wang2024physics] is used to demonstrate practical battery management. The dataset comprises experiments on 55 lithium-ion batteries ($\mathrm{LiNi_{0.5}Co_{0.2}Mn_{0.3}O_{2}}$), each with a nominal capacity of $2000\,\mathrm{mAh}$ and a nominal voltage of $3.6\,\mathrm{V}$. The experiments cover six charge and discharge regimes that emulate realistic operating conditions, including constant C rates, variable-rate profiles, randomized usage, and a satellite-use pattern.
The problem setting is to implement two key functions of the proposed system. The first task is to calibrate a coupled electrochemical–thermal model that can accurately simulate voltage and temperature dynamics under diverse cycling conditions. The second task is to develop predictive models for battery SOH estimation with quantified uncertainty, enabling risk-informed decision support for lifecycle management.
Realization of the Five-Tier System
We instantiate these tasks by mapping them to Tier 1–5 of the proposed system. Figure 2 illustrates how Tiers 1 and 2 are realized in the Omniverse-based 3D environment, where geometric models are bound with telemetry and simulation data to visualize battery states and predictions. We additionally leverage NVIDIA SimReady assets to accelerate the construction of standardized, simulation-ready battery components for Tiers 1 and 2, ensuring consistent geometric and physical representations within the Omniverse environment. Building on this foundation, Tier 3 is enabled through multi-physics modeling and calibration, while Tiers 4 and 5 leverage AI engines to provide predictive and prescriptive intelligence with uncertainty quantification.
Tier 3: Multi-physics Modeling and Calibration
@@
We integrate the SPM and DFN electrochemical models with a lumped thermal model using PyBaMM as the modeling platform [sulzer2021python]. The SPM offers computationally efficient simulations suitable for rapid, real-time predictions, while the DFN model provides detailed representation of lithium-ion transport processes and reaction kinetics. A lumped thermal model is coupled with the DFN to accurately capture thermal dynamics under diverse operating conditions.
@@
To ensure fidelity with respect to experimental battery performance, we apply Bayesian optimization for parameter calibration. This procedure adjusts over 15 model parameters (e.g., diffusivities, reaction rates, conductivities, electrode geometries, thermal conductivities) by minimizing discrepancies between simulated and measured voltage and temperature. Through iterative refinement, the calibrated models achieve strong consistency with experimental observations, forming the predictive backbone of the digital twin.
Tier 4: Physics-informed Neural Network
@@
Leveraging NVIDIA PhysicsNeMo, we construct PINNs trained on data generated from the calibrated multi-physics simulations. The loss function integrates both data-driven accuracy terms and physics-based constraints, penalizing violations of electrochemical principles and thermal equilibrium. By enforcing these constraints via automatic differentiation at strategically selected points, the PINN captures complex battery dynamics while maintaining physical consistency. This enables Tier 4 prescriptive functions by providing physically trustworthy predictions for SOH and RUL.
Tier 5: Uncertainty Aware Decision-Making
To support automatically decision-making in Tier 5, we incorporate Deep Autoencoding Gaussian Mixture Model (DAGMM) for uncertainty quantification. DAGMM combines an autoencoder with a Gaussian Mixture Model (GMM) to evaluate the likelihood of latent representations of operational data. The resulting energy-based score provides a principled uncertainty metric, identifying anomalies and quantifying confidence in predictions. By capturing uncertainty alongside predictions, this implementation equips the digital twin with the ability to support risk-informed management decisions.
Evaluation Results
@@
Rigorous evaluations are given to demonstrate the superior predictive accuracy and reliability of the proposed digital twin system through multi-physics model calibration and uncertainty quantifications.
(a) Voltage prediction under 1C, 2C, and 3C discharge rates.
(b) Temperature prediction under 1C, 2C, and 3C discharge rates.
@@
Figure 4: Multi-physics calibration results against the XJTU battery dataset after Bayesian optimization.
@@
For the multi-physics calibration, our model demonstrated good estimation accuracy compared with the XJTU battery dataset as shown in Figure 4. Across ten repeated trials, the calibrated model achieved an average voltage prediction MAPE of $0.92\pm 0.15\%$ (1C), $1.06\pm 0.17\%$ (2C), and $1.57\pm 0.21\%$ (3C). For temperature prediction, the MAPE reached $0.07\pm 0.01\%$ (1C), $0.18\pm 0.01\%$ (2C), and $0.39\pm 0.05\%$ (3C). These results validate the effectiveness of the multi-physics model in accurately reproducing voltage and temperature behaviors of batteries.
@@
Building upon the calibrated simulations, we can further assess the predictive performance of our PINN-based SOH prediction model, including its capability to quantify prediction uncertainty. As shown in Figure 5, our model demonstrated robust predictive accuracy in SOH, across ten runs achieving MAPE at $1.09\pm 0.04\%$. Most importantly, the integrated uncertainty quantification via the DAGMM provided reliable indications of prediction confidence. As shown in Figure 6, the energy-based uncertainty scores exhibited a strong positive correlation with the actual prediction errors, effectively highlighting scenarios where predictions were less reliable due to data distribution shifts.
@@
Future Directions
@@
In this section, we discuss several future directions that can be explored to further enhance the intelligent battery management using this digital twin architecture.
@@
Foundation Models for Battery Intelligence: Large Language Models (LLMs) are foundation models trained on massive text corpora using transformer architectures, enabling diverse tasks such as reasoning, summarization, and question answering without task-specific supervision. Beyond language, they serve as general-purpose engines for knowledge representation and code generation across scientific domains. Building on these capabilities, LLMs adapted on battery-related literature present promising opportunities for digital twins through automated knowledge extraction and model generation [zuo2025large]. Future implementations envision specialized battery foundation models that automatically design PINN architectures for specific chemistries, synthesize insights from vast research to identify degradation mechanisms, recommend experimental protocols, and support conversational interfaces for intuitive querying of battery states and explanations of complex degradation phenomena.
@@
Blockchain-based Battery Passport Systems: The implementation of blockchain in battery lifecycle management offers transformative potential for future battery passport systems, enabling complete traceability and transparency. Distributed ledger technologies will record manufacturing data, operational history, maintenance, and performance metrics as immutable entries, ensuring reliable information sharing across manufacturers, operators, and recyclers [dasgupta2025intelligent]. Smart contracts could automate management decisions such as maintenance scheduling and end-of-life processing, reducing fraud in condition reporting. Blockchain-based passports will further support circular economy initiatives by providing verified health data for second-life applications, allowing batteries retired from vehicles to be reused in stationary storage with confidence.
@@
Differentiable Simulation for Optimal Control: The development of fully differentiable battery simulation frameworks enables gradient-based optimization of battery operational strategies directly through physics-based models, eliminating the need for computationally expensive reinforcement learning approaches. Future implementations could leverage automatic differentiation through simulation platforms to enable direct optimization of charging protocols, thermal management strategies, and load balancing decisions with respect to battery health and performance objectives. This approach provides enhanced transparency compared to black-box optimization methods while enabling principled handling of operational constraints through physics-based penalty terms in the optimization objective.
Conclusion
@@
To promote the transition of BMS towards intelligent systems for achieving full-lifecycle efficient management across diverse energy storage applications, we presented an intelligent battery management concept based on a five-tier digital twin system for autonomous operations. We first highlighted current BMS challenges and the need for advanced digital twin approaches. Then, we introduced the system design, progressing from geometric modeling to autonomous operation, supported by PIML and uncertainty quantification. Its applications span predictive health management, operational optimization, and lifecycle control. Finally, a system implementation validated our design, showing sub-1% voltage/temperature errors and robust SOH predictions with MAPE below 3%, demonstrating the system’s potential for interpretable modeling and reliable decision-making.